# NAAQA: A Neural Architecture for Acoustic Question Answering

### v1

@@ 4 Method
Our goal is to devise a system that takes as inputs a variable length text string representing a question and a variable duration audio signal representing an acoustic scene. The system’s task is to learn to interpret the question in the context of the acoustic scene and to generate an answer. We first describe the Visual FiLM architecture [undefab] that is used as a baseline, then the NAAQA model and finally how coordinate maps[undefaam] can be applied in an acoustic context. Both the NAAQA and Visual FiLM share an overall common architecture which is depicted in Figure 2.
@@ 4.1 Baseline model: Visual FiLM
The FiLM architecture [undefab] is inspired by Conditional Batch Normalization architectures [undefaan]. It achieved state of the art results on the CLEVR [undeff] VQA task at the time of publication. We refer to this neural network as Visual FiLM or baseline model in subsequent sections. The network takes a visual scene and a text-based question as inputs and predicts an answer to the question for the given scene. The question goes through the text-processing module which uses $G$ unidirectional gated recurrent units (GRUs) to extract context from the text input (yellow area in Figure 2). The visual scene is processed by the convolutional module (blue area in the figure). The first step of this module is feature extraction (orange box), performed by a Resnet101 model [undefaae] pre-trained on ImageNet [undefaao]. The extracted features are processed by a convolutional layer with batch normalization [undefaan] and ReLU [undefaap] activation followed by $J$ Resblocks illustrated in details in the red area in the figure. Unless otherwise specified, batch normalization and ReLU activation functions are applied to all convolutional layers. Each Resblock $j$ comprises convolutional layers with $M$ filters that are linearly modulated by FiLM layers through the two $M\times 1$ vectors $\boldsymbol{\beta}_{j}$ (additive) and $\boldsymbol{\gamma}_{j}$ (multiplicative). This modulation emphasizes the most important feature maps and inhibits the irrelevant maps given the context of the question. $\boldsymbol{\beta}_{j}$ and $\boldsymbol{\gamma}_{j}$ are learned via fully connected layers using the text embeddings extracted by the text processing module as inputs (purple area in the figure). The affine transformation in the batch normalization before the FiLM layer is deactivated. The FiLM layer applies its own affine transformation using the learned $\boldsymbol{\beta}_{j}$ and $\boldsymbol{\gamma}_{j}$ to modulate features. Several Resblocks can be stacked to increase the depth of the model, as illustrated in Figure 2. Finally, the classifier module is composed of a $1\times 1$ convolutional layer [undefaaq] with $C$ filters followed by max pooling and a fully connected layer with $H$ hidden units and an output size $O$ equal to the number of possible answers (Gray in Figure 2). A softmax layer predicts the probabilities of the answers. In order to use the Visual FiLM as a baseline for our experiments, we extract a 2D spectro-temporal representation of the acoustic scenes as depicted at the bottom of Figure 2. This is fed to the model as if it was an image which is the simplest way to adapt the unmodified visual architecture to acoustic data.
@@
Figure 2: Common Architecture. Two inputs: a spectro-temporal representation of an acoustic scene and a textual question. The spectro-temporal representation goes through a feature extractor (Parallel feature extractor detailed in Section 4.2 for NAAQA and Resnet101 pretrained on ImageNet for Visual FiLM) and then a serie of $J$ Resblocks that are linearly modulated by $\boldsymbol{\beta}_{j}$ and $\boldsymbol{\gamma}_{j}$ (learned from the question input) via FiLM layers. Coordinate maps are inserted before convolution operation with a pink border. The output is a probability distribution of the possible answers.
@@
4.2 NAAQA: Adapting feature extractors to acoustic inputs
As in Visual FiLM, the first step in the NAAQA model is feature extraction (orange box in Figure 2). Because acoustic signals present unique properties, two feature extraction modules that are specifically tailored to modeling sounds are introduced :
  • the Parallel feature extractor (Figure 3a) processes time and frequency features independently in parallel pipelines.
  • the Interleaved feature extractor (Figure 3b) captures time and frequency features in a single convolutional pipeline.
@@ 4.2 NAAQA: Adapting feature extractors to acoustic inputs
In both cases, the feature extractor is trained end-to-end with the rest of the network and uses a combination of 1D convolutional filters to process a 2D spectro-temporal representation. The 1D filters process the time and frequency axis independently as opposed to the 2D filters typically used in image processing.
@@
The Parallel feature extractor (Figure 3(a)) captures time and frequency features in parallel pipelines. The frequency pipeline (green in the figure) comprises a serie of $K$ frequency blocks. Each block is composed of a 1D convolution with $N_{K}$ $3\times 1$ kernels and $2\times 1$ strides followed by a $1\times 2$ maxpooling. With a stride larger than $1\times 1$, the convolution operation downsamples the frequency axis and the pooling operation downsamples the time axis. This downsampling strategy allows features in both parallel pipelines to be of the same dimensions. The time pipeline (yellow in the figure) is the same as the frequency pipeline except that the convolutional kernel operates along the time dimension and the pooling along the frequency dimension. The convolution kernel is $1\times 3$ and the pooling kernel is $2\times 1$. The activation maps of both pipelines are concatenated channel-wise and a representation combining both the time and frequency features is created using a $1\times 1$ convolution [undefaaq] with $P$ filters and a stride of one. The feature maps dimensionality is either compressed or expanded depending on the number of filters $P$ in the $1\times 1$ convolution. The $1\times 1$ convolution can also be removed thus leaving it up to the next $3\times 3$ convolution to fuse the time and frequency features.
@@
The Interleaved feature extractor (Figure 3(b)) processes the input spectrogram in a single pipeline composed of a serie of $K$ interleaved blocks (purple in the figure). Each block comprises a $1\times 3$ time convolution with $N_{K}$ filters and stride $1\times 2$ followed by a $3\times 1$ frequency convolution with $N_{K}$ filters and stride $2\times 1$. A $1\times 1$ convolution with $P$ filters processes the output of the last block to either compress or expand its dimensionality. As an alternative configuration, the order of the convolution operation in each block can be reversed so that it first operate along the frequency axis and then the time axis. Compared to the Parallel feature extractor, learned time-frequency representations are created after each block instead of only at the end of the pipeline.
@@
For both extractors, the convolutions in the first block comprise $N_{1}$ convolutional filters and the number of filters is doubled after each block. More blocks $K$ results in more down sampling of the feature maps shape which brings down the computational cost of the model.
4.3 Coordinate maps with acoustic inputs
The baseline model makes use of the CoordConv [undefaam] technique when tackling the Visual Question Answering task (pink border boxes in Figure 2). This technique provides a reference coordinate system that is consistent across all input features. It comprises 2 feature maps filled with (constant, untrained) coordinate information that are concatenated channel-wise to the input features of a convolutional layer. The first concatenated map is a 2D matrix where all values for a given column are the same. The values start from -1 on the first row and increase linearly up to 1 on to last row. The second map is equivalent to the first except that the linearly increasing axis is the column and the fixed axis is the row. These coordinate maps help the convolutional layer learn localization features, at least in the visual domain. Coordinate maps can be inserted before any convolution operation at any depth of the neural network.
@@ 4.3 Coordinate maps with acoustic inputs
In the visual domain both axis of an image encode spatial information. With spectro-temporal representations, the Y axis corresponds to frequency and the X axis corresponds to time. In this context, we call the first map frequency coordinate map and the second map time coordinate map. [undefaar] [undefaar] proposed Frequency-Aware convolutions which is equivalent to concatenating only a frequency coordinate map to the input of a convolution layer. In this work, the performance of both Frequency and Time coordinate maps are investigated. All spectro-temporal representations in the CLEAR dataset have the same range for the frequency axis but the range for the time axis varies depending on the duration of the acoustic scenes. Also, there are questions directly related to time localization whereas frequency localization is more related to pattern recognition. We hypothesize that time coordinate maps might have a larger impact on performance because they provide a relative time scale that the model can use to enhance its temporal localization capabilities.
5 Experiments & Results
We first describe the preprocessing of the acoustic signals and the experimental conditions. We then study the impact of different feature extractors on the CLEAR task. We examine how the insertion of coordinate maps at different depths in the architecture affects performance. Finally, we do an ablation study to minimize the complexity of NAAQA and we study the impact of the dataset size on the training process.
5.1 Acoustic Processing
The raw acoustic signal sampled at 48 kHz is transformed to create a 2D time-frequency representation where the frequency axis is log-scaled. After experimenting with different window sizes and time shifts, we found that spectrograms using a Hanning window of 512 samples ($\sim$106 msec) and a time shift of 2048 samples ($\sim$427 msec) gave the best results. Using a time shift larger than the window create a spectro-temporal representation for which the time axis has been downsampled. This provides a cost-effective representation, given that the elementary sounds used to create the scenes are slowly varying in time. We refer to this downsampled spectrogram as spectrogram in the remaining parts of the paper. A 512 points FFT is used yielding 256 frequency bins per window. $N$ spectra are concatenated, where $N$ is equal to the number of windows. Since the duration (and therefore $N$) is not constant, we pad the spectrogram time axis with zeros so that all spectrograms have the same dimensions. The power spectrum is normalized using the mean and standard deviation of the training data. This normalization scheme brings the mean close to 0 and the standard deviation close to 1 which has been shown to speed up convergence [undefaas].
@@ 5.2 Experimental conditions
The networks presented in subsequent sections are trained on the CLEAR dataset which comprises 50 000 scenes and 4 questions per scene for a total of 200 000 records from which 140 000 (70%) are used for training, 30 000 (15%) for validation and 30 000 (15%) for test. All models were trained for a maximum of 40 epochs or until the validation loss stops decreasing for 6 consecutive epochs. The Adam optimizer [undefaat] with a learning rate of $3\text{\times}{10}^{-4}$ was used to optimize the cross entropy loss. L2 regularization with a weight decay of $5\text{\times}{10}^{-6}$, dropout [undefaau] with a drop probability of 0.25 and batch normalization [undefaan] are used to regularize the learning process. A batch size of 128 is used for all experiments. The results are reported in terms of accuracy, that is the percentage of correct answers on the total. Since initialization of deep architectures has a profound impact on training convergence, we developed a python library torch-reproducible-block [^2] The link will be made available upon publication to control the model initial conditions and design reproducible experiments. To ensure the robustness of the results, each model is trained 5 times with 5 different random seeds [^3] The link will be made available upon publication. To reduce the effect of getting stuck in local minima, the accuracies presented in subsequent sections are the average of the 3 best runs.
5.3 Initial configuration
@@
The initial configuration for the proposed model comprises $G=4096$ GRU units, $J=4$ Resblocks with $M=128$ filters each and a classifier composed of a $1\times 1$ convolution with $C=512$ filters and $H=1024$ hidden units in the fully connected layer. This configuration includes both time and frequency coordinate maps in each location highlighted in pink in Figure 2. The above hyper-parameters are also used for the baseline model described in Section 4.1.
5.4 Feature Extraction
Accuracy by question type on the test set(%)
Extractor Absolute Global Relative Count Count
Type Overall Instrument Note Brightness Loudness Exist Position Position Position Count Comparison Instruments
Resnet101 (Baseline) 77.9 ±0.14 85.0 ±0.56 67.7 ±1.54 93.8 ±0.20 93.6 ±0.46 81.6 ±0.46 74.4 ±1.83 83.8 ±3.76 56.4 ±1.70 51.2 ±0.66 64.1 ±0.64 37.9 ±0.62
Parallel (Fig 3(a)) 89.8 ±0.30 95.6 ±0.59 92.3 ±1.00 97.9 ±0.05 97.7 ±0.24 87.3 ±0.28 95.3 ±0.06 96.5 ±0.12 59.4 ±2.79 62.9 ±1.01 63.0 ±0.62 38.1 ±3.17
Interleaved Freq (Fig 3(b)) 59.6 ±12.47 48.0 ±19.78 34.2 ±20.30 82.0 ±6.59 80.3 ±8.13 70.7 ±6.16 51.9 ±20.78 73.9 ±11.36 43.9 ±4.99 40.1 ±4.50 55.6 ±1.53 37.9 ±2.83
Interleaved Time (Fig 3(b)) 90.9 ±0.21 96.8 ±0.19 93.8 ±0.47 98.6 ±0.17 98.2 ±0.21 88.6 ±0.24 96.4 ±0.53 96.9 ±0.56 70.5 ±1.40 64.6 ±0.80 64.0 ±0.39 41.4 ±1.89
TABLE III: Impact of different feature extractors. The initial configuration is used for the rest of the network (defined in Section 5.3). Test accuracy is reported by question type (Examples for each question type are available in Table I).
We investigated different feature extraction strategies with the initial configuration. Results are presented in Table III.
The baseline uses a Resnet101 module [undefaae] pre-trained on Imagenet [undefaao] as feature extractor. This extractor is expecting a 3 channels input but the spectro-temporal representation of the acoustic scene comprise only 1 channel. To work around this constraint, the amplitude of the frequency bins of the spectrogram are transformed into RGB values using the viridis matplotlib colormap [^4] https://matplotlib.org/examples/color/colormaps_reference.html thus creating a 3 channels input Spectrograms have a very different structure than visual scenes and, as expected, the pre-learnt knowledge gathered in a visual context does not transfer directly to the acoustic context. This configuration does however achieve a surprisingly good overall accuracy of 77.9%. It performs quite well with questions related to brightness and loudness with 93.8% and 93.6% respectively. One possible explanation is that both brightness and loudness are directly related to the distribution of colors along the frequency axis in the spectrograms. These features are efficiently encoded by the visual representations learned by the Resnet101 model. The baseline does not perform as well with questions related to absolute position which is somewhat surprising because these questions correspond to a similar pattern matching problem on the time axis instead of the frequency axis. It does however perform better with global position questions. This is logical because this type of question refer to an approximate position (beginning, middle, end) which constrain the number of possible answers to 3 instead of 15. The baseline model has difficulties with questions related to notes (67.7%). This might be explained by the fact that a note can be identified by its fundamental frequency and harmonics which can be far apart on the frequency axis. Visual models are trained to recognize localized features and therefore struggle to recognize notes. This is a feature that is typical of acoustic signals.
The Parallel feature extractor (Figure 3(a)) which processes 2D spectrogram ($256\times\text{N}\times 1$) using independent time and frequency pipelines reaches an overall accuracy of 89.8%. It performs well on all questions types except relative position and count. These results show that building complex time and frequency feature separately and fusing them at a later stage is a good strategy to learn acoustic features for this task.

### v2

@@ 4 Method
We first describe the original Visual FiLM architecture [undefz] that we use as baseline model, then the proposed modifications that lead to our NAAQA architecture and, finally, NAAQA with a MALiMo module.
@@ 4.1 Baseline model: Visual FiLM
Both the proposed NAAQA and Visual FiLM [undefz] share an overall common architecture which is depicted in Figure 2. Visual FiLM, that we will use as baseline model, is inspired by Conditional Batch Normalization architectures [undefaap] and achieved state of the art results on the CLEVR VQA task [undeff]. The network takes a visual scene and a text-based question as inputs and predicts an answer to the question for the given scene. The text-processing module uses $G$ unidirectional gated recurrent units (GRUs) to extract context from the text input (yellow area in Figure 2). The visual scene is processed by the convolutional module (blue area in the figure). The first step of this module is feature extraction (orange box), performed by a Resnet101 model [undefaae] pre-trained on ImageNet [undefaaq]. The extracted features are processed by a convolutional layer with batch normalization [undefaap] and ReLU [undefaar] activation followed by $J$ Resblocks illustrated in details in the red area in the figure. Unless otherwise specified, batch normalization and ReLU activation functions are applied to all convolutional layers. Each Resblock $j$ comprises convolutional layers with $M$ filters that are linearly modulated by FiLM layers through the two $M\times 1$ vectors $\boldsymbol{\beta}_{j}$ (additive) and $\boldsymbol{\gamma}_{j}$ (multiplicative). This modulation emphasizes the most important feature maps and inhibits the irrelevant maps given the context of the question. $\boldsymbol{\beta}_{j}$ and $\boldsymbol{\gamma}_{j}$ are learned via fully connected layers using the text embeddings extracted by the text processing module as inputs (purple area in the figure). The affine transformation in the batch normalization before the FiLM layer is deactivated. The FiLM layer applies its own affine transformation using the learned $\boldsymbol{\beta}_{j}$ and $\boldsymbol{\gamma}_{j}$ to modulate features. Several Resblocks can be stacked to increase the depth of the model, as illustrated in Figure 2. Finally, the classifier module is composed of a $1\times 1$ convolutional layer [undefaas] with $C$ filters followed by max pooling and a fully connected layer with $H$ hidden units and an output size $O$ equal to the number of possible answers (Gray in Figure 2). A softmax layer predicts the probabilities of the answers. In order to use the Visual FiLM as a baseline for our experiments, we extract a 2D spectro-temporal representation of the acoustic scenes as depicted at the bottom of Figure 2. The Resnet101 pre-trained extractor expects a 3 channels visual input but the spectro-temporal representation comprises only 1 channel. To work around this constraint, the spectro-temporal information is simply repeated 3 times thus creating a 3 channels input (only when using Resnet101 as feature extractor). This modified spectro-temporal representation is then fed to the model as if it was an image which is the simplest way to adapt the unmodified visual architecture to acoustic data. We call this architecture Visual FiLM Resnet101.
@@
Figure 2: Common Architecture. Two inputs: a spectro-temporal representation of an acoustic scene and a textual question. The spectro-temporal representation goes through a feature extractor (Parallel and Interleaved feature extractor detailed in Section 4.2.1 for NAAQA and Resnet101 pretrained on ImageNet for Visual FiLM) and then a serie of $J$ Resblocks that are linearly modulated by $\boldsymbol{\beta}_{j}$ and $\boldsymbol{\gamma}_{j}$ (learned from the question input) via FiLM layers. Coordinate maps are inserted before convolution blocks with a pink border. The output is a probability distribution of the possible answers.
@@
4.2 The proposed NAAQA architecture
In order to define the NAAQA architecture, we performed a number of modifications to the baseline architecture that will be described in the following sections.
4.2.1 Feature Extraction
As in Visual FiLM, the first step in the NAAQA model is feature extraction (orange box in Figure 2). The most obvious adaptation of Visual FiLM to acoustic data is to retrain the feature extraction module on the scenes from CLEAR2. In order to do this, we used three 2D convolutional layers, with $3\times 3$ kernels, stride $2\times 2$ and $N_{1}$, $N_{2}$, and $N_{3}$ filters respectively followed by a $1\times 1$ convolutions with $N_{4}$ filters. We will refer to this model as NAAQA 2D Conv.
@@ 4.2.1 Feature Extraction
However, because acoustic signals present unique properties, we introduce two feature extraction modules that are specifically tailored to modeling sounds: the Parallel feature extractor (Figure 3a) processes time and frequency features independently in parallel pipelines; the Interleaved feature extractor (Figure 3b) captures time and frequency features in a single convolutional pipeline. In both cases, the feature extractor is trained end-to-end with the rest of the network and uses a combination of 1D convolutional filters to process a 2D spectro-temporal representation. The 1D filters process the time and frequency axis independently as opposed to the 2D filters typically used in image processing.
@@
The design of the Parallel feature extractor (Figure 3(a)) is inspired by the work of [undefaai] [undefaai] where 1D filters are used to capture time and frequency features separately. While [undefaai] time-frequency model includes only 1 time and 1 frequency convolution which are then concatenated together, our extractor stacks multiple 1D convolutions in two parallel pipelines. The time and frequency features are only fused at the end of both pipelines. This yields more complex features. The frequency pipeline (green in the figure) comprises a serie of $K$ frequency blocks. Each block is composed of a 1D convolution with $N_{K}$ $3\times 1$ kernels and $2\times 1$ strides followed by a $1\times 2$ maxpooling. With a stride larger than $1\times 1$, the convolution operation downsamples the frequency axis and the pooling operation downsamples the time axis. This downsampling strategy allows features in both parallel pipelines to be of the same dimensions. The time pipeline (yellow in the figure) is the same as the frequency pipeline except that the convolutional kernel operates along the time dimension and the pooling along the frequency dimension. The convolution kernel is $1\times 3$ and the pooling kernel is $2\times 1$. The activation maps of both pipelines are concatenated channel-wise and a representation combining both the time and frequency features is created using a $1\times 1$ convolution [undefaas] with $P$ filters and a stride of one. The feature maps dimensionality is either compressed or expanded depending on the number of filters $P$ in the $1\times 1$ convolution. We call the corresponding model NAAQA Parallel. The $1\times 1$ convolution can also be removed thus leaving it up to the next $3\times 3$ convolution to fuse the time and frequency features.
@@
The Interleaved feature extractor (Figure 3(b)) processes the input spectrogram in a single pipeline composed of a serie of $K$ interleaved blocks (purple in the figure). Each block comprises a $1\times 3$ time convolution with $N_{K}$ filters and stride $1\times 2$ followed by a $3\times 1$ frequency convolution with $N_{K}$ filters and stride $2\times 1$. A $1\times 1$ convolution with $P$ filters processes the output of the last block to either compress or expand its dimensionality. We call the corresponding model NAAQA Interleaved Time.
As an alternative configuration, the order of the convolution operation in each block can be reversed so that it first operates along the frequency axis and then the time axis. The model is called NAAQA Interleaved Freq, in this case. Compared to the Parallel feature extractor, time-frequency representations are created after each block instead of only at the end of the pipeline.
@@
For all extractors, the convolutions in the first block comprise $N_{1}$ convolutional filters and the number of filters is doubled after each block ($N_{i}$ = $2N_{i-1}$). More blocks (higher $K$) results in more down sampling of the feature maps shape which brings down the computational cost of the model.
4.2.2 Coordinate maps for acoustic inputs
The Visual FiLM model concatenates coordinate maps (CoordConv [undefaaj]) to the input of convolutional layers (pink border boxes in Figure 2) when tackling the VQA task. In the visual domain both axis of an image encode spatial information. Coordinate maps have, therefore, the same meaning in the $x$ or $y$-axis.
@@ 4.2.2 Coordinate maps for acoustic inputs
In spectro-temporal representations for audio, however, the $y$-axis corresponds to frequency and the $x$-axis to time. We, therefore, call the maps frequency and time coordinate map, respectively. All spectro-temporal representations in CLEAR2 have the same range for the frequency axis but the range for the time axis varies depending on the duration of the acoustic scenes. We hypothesize that time coordinate maps might have a stronger impact on performance because they provide a relative time scale that the model can use to enhance its temporal localization capabilities.
4.2.3 Complexity Optimization
We performed optimization of the most important hyper-parameters in the NAAQA architecture with the goal of reducing model complexity. These include number of GRU text-processing units $G$; the number of Resblock $J$ that dictates the number of FiLM layers and, therefore, the number of modulation coefficients to compute; the number of convolutional filters $M$ in each block; the number of filters $C$ and the number of hidden units $H$ in the classifier module. We refer to the resulting model by prepending Optimized to the model name.
4.2.4 NAAQA with a MALiMo module
In MALiMo [undefaa], [undefaa] add a second set of FiLM layers that acts has an auxiliary controller. The controller uses the extracted acoustic features to further modulate the intermediate Resblocks. To evaluate the impact of MALiMo on CLEAR2, a MALiMo module was added to NAAQA. We refer to this configuration by appending MALiMo ctrl to the names introduced above. In our implementation of the module we replaced LSTMs with GRUs and adapted the inputs to the acoustic features that we study.
5 Experiments
We perform experiments to compare the effect on performance of our modification to the baseline model. Most of our experiments are performed on the proposed CLEAR2 data. We first investigate the performance of different feature extraction methods and compare them to the Visual FiLM Resnet101 feature extractor. Then, we show the effect of time and frequency coordinate maps at different levels of the model. Moreover, we perform an hyper-parameters ablation study to reduce the complexity of the model. We finally test the addition of a MALiMo module to our model. In order to demonstrate the generality of our results, we compare the performance of our model on the CLEAR2 and DAQA datasets.
5.1 Acoustic Pre-processing
The raw acoustic signal (sampled at 48 kHz for CLEAR2 and 16kHz for DAQA) is processed to create a 2D time-frequency representation using Mel scale [undefaat] spectrograms. After some experimentation we decided to use 64 Mel coefficients for both CLEAR2 and DAQA computed over samples weighted by a Hanning window. For CLEAR2 the window size was of 512 samples ($\sim$10.6 msec) whereas for DAQA we used windows of 400 samples ($\sim$25ms) as in [undefaa]. The time shift between consecutive windows (stride) was also optimized depending on the properties of the audio data. For CLEAR2 we found the best results for a time shift of 2048 samples ($\sim$42.7 msec). This is possible because of the properties of the sustained notes in the data set which vary slowly in time. Using such a long time shift allowed us to reduce more than ten folds the computational costs in our experiments. Because the DAQA dataset contains sounds that are more quickly varying, the same optimization is not possible. We tested with time shifts of 160 samples (10ms) and 1600 samples (100ms) observing a 5% drop in accuracy for the longer stride. All results based on CLEAR2 are reported with long window shift (long stride), with the exception of the comparison between short and long stride on both CLEAR2 and DAQA in Supplementary C.
Because the duration of the scenes in CLEAR2 is not constant, we pad the spectrogram along the time axis with zeros so that all spectrograms have the same dimensions ($1\times 64\times 418$) which corresponds to a maximum length of $\sim$17.9 sec. The power spectrum is normalized to the mean and standard deviation of the training data with the goal of speeding up convergence [undefaau].
@@ 5.2 Experimental conditions
Unless specified otherwise, the models presented in subsequent sections are trained on the CLEAR2 dataset which comprises 50 000 scenes and 4 questions per scene for a total of 200 000 records from which 140 000 (70%) are used for training, 30 000 (15%) for validation and 30 000 (15%) for test. The test set is generated using a different set of elementary sound which ensure that the network can not memorize the elementary sounds and therefore acts has a better benchmark for generalization. Optimisation technique and other training settings are further described in Supplementary A. The results are reported in terms of accuracy, that is the percentage of correct answers on the total. Since initialization of deep architectures has a profound impact on training convergence, we developed a python library torch-reproducible-block [^3] The link will be made available upon publication to control the model initial conditions and design reproducible experiments. To ensure the robustness of the results, each model is trained 5 times with 5 different random seeds.
5.3 Initial model configuration
@@
The initial configuration for the proposed model comprises $G=4096$ GRU units, $J=4$ Resblocks with $M=128$ filters each and a classifier composed of a $1\times 1$ convolution with $C=512$ filters and $H=1024$ hidden units in the fully connected layer. This configuration includes both time and frequency coordinate maps in each location highlighted in pink in Figure 2.
Number of Overall Accuracy by question type (%)
Configuration Parameters Acc. Instrument Note Brightness Loudness Exist Abs. Pos. Glob. Pos. Rel. Pos. Count Count Comp. Count Inst.
Baselines
Random Answer 1.75 1.75 1.75 1.75 1.75 1.75 1.75 1.75 1.75 1.75 1.75 1.75
Most common Answer (Yes) 7.3 0 0 0 0 55.62 0 0 0 0 47.27 0
Visual FiLM Resnet101 6.71 M 62.3 ±0.78 61.9 ±0.85 37.8 ±1.30 83.9 ±0.58 81.9 ±0.61 73.2 ±0.70 51.7 ±1.40 73.1 ±2.29 48.4 ±1.82 41.5 ±0.42 59.8 ±0.57 39.9 ±3.03
NAAQA
NAAQA 2D Conv 5.61 M 77.6 ±0.72 80.6 ±0.66 73.4 ±1.28 90.5 ±1.44 86.5 ±0.76 81.3 ±0.57 74.7 ±2.50 87.5 ±0.42 54.0 ±1.43 53.8 ±1.01 60.8 ±1.34 49.0 ±3.53
NAAQA Interleaved Freq 5.61 M 67.2 ±0.98 63.0 ±1.93 48.0 ±2.14 84.5 ±1.07 81.1 ±0.52 72.1 ±1.24 65.1 ±1.04 80.1 ±0.66 49.3 ±1.59 42.5 ±1.51 56.6 ±2.45 46.5 ±1.57
NAAQA Interleaved Time 5.61 M 78.0 ±0.51 81.8 ±1.06 70.6 ±1.24 90.9 ±0.78 87.9 ±0.94 81.6 ±0.66 75.9 ±1.08 87.4 ±0.25 60.0 ±4.07 53.6 ±0.63 61.2 ±0.73 50.0 ±2.23
NAAQA Parallel 5.61 M 78.5 ±0.45 80.4 ±0.83 72.6 ±1.19 91.2 ±0.74 87.1 ±0.28 80.4 ±0.16 78.7 ±1.26 88.8 ±0.53 55.4 ±1.63 52.6 ±0.43 59.8 ±2.13 50.2 ±3.49
Optimized NAAQA Parallel 1.68 M 79.5 ±0.05 81.7 ±0.20 74.2 ±0.58 91.9 ±0.17 87.4 ±0.31 81.2 ±0.10 79.3 ±0.77 90.0 ±0.15 58.0 ±1.32 53.8 ±1.40 60.6 ±0.89 50.4 ±1.10
NAAQA + MALiMo
Visual FiLM Resnet101 + MALiMo ctrl 6.82 M 63.6 ±1.20 61.8 ±1.40 36.3 ±1.41 83.5 ±0.86 82.3 ±0.59 73.4 ±1.06 57.1 ±3.58 76.1 ±1.94 47.4 ±2.98 41.6 ±0.80 56.7 ±4.23 43.2 ±3.64
NAAQA 2D Conv + MALiMo ctrl 6.71 M 77.1 ±1.12 79.2 ±1.01 71.3 ±1.54 90.5 ±0.81 86.1 ±0.73 80.3 ±0.77 75.7 ±2.08 87.8 ±0.74 52.7 ±1.94 51.9 ±2.73 59.0 ±3.00 50.2 ±2.48
NAAQA Interleaved Freq + MALiMo ctrl 6.71 M 64.8 ±3.91 55.3 ±8.33 43.1 ±8.77 83.2 ±1.33 80.0 ±1.75 70.3 ±2.74 63.6 ±4.72 78.7 ±3.27 45.8 ±3.53 41.4 ±1.96 55.1 ±2.75 48.7 ±2.05
NAAQA Interleaved Time + MALiMo ctrl 6.71 M 77.1 ±0.63 79.9 ±0.86 70.8 ±1.10 90.5 ±1.00 86.7 ±0.89 80.3 ±0.34 74.5 ±1.06 87.6 ±0.35 55.2 ±2.39 53.8 ±0.43 59.9 ±2.32 50.0 ±0.93
NAAQA Parallel + MALiMo ctrl 6.71 M 77.3 ±0.93 78.2 ±1.50 71.3 ±1.00 89.9 ±0.66 85.8 ±0.67 79.7 ±0.45 77.7 ±2.02 87.6 ±0.53 53.6 ±1.84 51.9 ±1.82 59.0 ±2.29 48.1 ±1.65
Optimized NAAQA Parallel + MALiMo ctrl 2.78 M 78.2 ±0.06 80.8 ±0.41 72.4 ±0.33 89.6 ±0.62 86.2 ±0.04 79.7 ±0.10 79.3 ±0.22 88.4 ±0.36 54.0 ±0.33 52.2 ±0.70 58.8 ±0.45 45.6 ±0.00
TABLE III: Results on CLEAR2. The table shows number of parameters, average accuracy (%), and standard deviation over 5 repetitions of the training. Overall accuracy as well as question-kind dependent accuracy are reported. The different configurations are reported in the same order as they are presented in the paper. The most common answer is “Yes”.
Configuration # Parameters Train Acc. Val Acc. Test Acc. Trainig time
Optimized NAAQA 2D Conv 1.68 M 65.2 ±0.64 58.0 ±0.82 58.3 ±0.98 0 days 06:48:12
Optimized NAAQA Parallel 1.68 M 66.6 ±0.51 60.4 ±0.08 60.4 ±0.21 0 days 06:49:52
Optimized NAAQA 2D Conv + MALiMo ctrl 2.78 M 58.5 ±3.03 54.2 ±2.42 54.4 ±2.35 0 days 07:31:37
Optimized NAAQA Parallel + MALiMo ctrl 2.78 M 67.3 ±1.22 64.1 ±0.54 64.3 ±0.72 1 days 04:59:46
TABLE IV: Results on DAQA The table presents number of parameters, average training, validation and test accuracy (%) with standard deviation over 5 repetitions of the training as well as average training time. Results are reported for four configurations, with and without the MALiMo module in the same order as they are presented in the paper.
6 Results and Discussion
Main results on the CLEAR2 data set are reported in Table III. It reports the complexity of the models in terms of number of parameters, the overall accuracy and accuracy dependent on question’s type. First we report two theoretical baselines given by random chance and majority class answers. Then we report results based on the baseline model Visual FiLM Resnet101 using the initial configuration described in Section 5.3. This architecture achieves the lowest accuracy of all tested models with 62.3%. This suggests that Mel spectrograms have a very different structure than visual scenes and, as expected, the pre-learned knowledge gathered in a visual context does not transfer directly to the acoustic context.
6.1 NAAQA modifications
Unless specified otherwise, the initial configuration described in section 5.3 is used for all models in this section.
6.1.1 Feature Extraction
The first improvement to the baseline is given by introducing a specific audio feature extraction module based on 2D convolutions. The NAAQA 2D Conv model has slightly fewer parameters than Visual FiLM Resnet101 because of the simpler feature extraction module and a much higher overall accuracy of 77.6%.